OpenStudy (anonymous):

D = R x T Two cars leave town at the same time going in the same direction. One travels 55 mi/h and the other travels 48 mi/h. In how many hours will they be 49 miles apart?

OpenStudy (anonymous):

OK so car 1 is modeled by D=55x and car 2 D=48x when can then model the difference in distance by saying that 49= 55x - 48x then solving for x which you will find to be 7 h

OpenStudy (anonymous):

ok, any other steps?

OpenStudy (anonymous):

well to solve it looks like 7x=49 x=49/7 x=7

OpenStudy (anonymous):

x represents time in hours

OpenStudy (anonymous):

ok tyvm, would you mind answering 1 more question for me? I think im starting to get it.

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

An airplane flew for 5 hours with a tail wind of 25 km/h. The return flight against the same wind took 5.5 hours. Find the speed of the airplane in still air.

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

the flight there would be modeled as D= 5h(V) + 5h(25km/h) and the flight back would be D=5.5h(V) - 5.5h(25km/h) since the distances are the same and the (v) which is velocity is the same we can equate the 2 equations so we get 5.5h(v) -137.5 = 5(v) + 125 then solve for (v) which is the speed of the plane in both directions :P I got 525km/h

OpenStudy (anonymous):

can you stretch out the steps a little more please?

OpenStudy (anonymous):

the plane trip there takes 5h SO 5h times velocity + 5 hours times wind speed = distance the reverse trip will be 5.5h times velocity - 5.5 hours times wind speed cause its against the wind this time. because distances are the same IE the plane travels the same distance both ways we can sub the first equation into the second for where D is as the first equation is = to D then we solve for velocity 5.5h(v) -137.5 = 5(v) + 125 5(v) - 5.5(v) = -137.5-125 -0.5(v) = -262.5 (v) = -262.5/-0.5 (v) = 525km/h

OpenStudy (anonymous):

Thank you so very much I get it now!

OpenStudy (anonymous):

no problem! good luck!